find the sum of the arithmetic series.-4,+1+6+11+........+91
Answer by stanbon(75887)

find the sum of the arithmetic series.-4,+1+6+11+........+91
----------------
a(1) = -4
d = 5
---
91 = -4 + (n-1)5
95 = (n-1)5
n-1 = 19
n = 20
---
S(20) = (20/2)(-4+91)
S(20) = 10(95)
S(20) = 950
=================
Cheers,
Stan H.